The physician ordered: 15mEq of potassium chloride. Available: 5mEq/10ml. How many teaspoons should be given?
Answer #1

ANSWER.

* Calculation of teaspoon of drug required.

     Formula

           Volume(ml) = Desired dose(D)/Dose in hand(H) * Quantity(Q).

Here

       Desired dose(D) = 15 mEq.

       Dose in hand(H) = 5 mEq.

          Quantity(Q) = 10 ml.

As per above formula

     Volume(ml) of drug required = D/H * Q

                                                   = 15/5 * 10

                                                   = 3 * 10

                                                   = 30 ml.

                 1 teaspoon = 5 ml

      Teaspoon of drug required = 30/5

                                                 = 6 teaspoon